OpenStudy (anonymous):

Solve the system of equations. 3x - 4y = -10 and 5x + 8y = -2 A. (1, 2) B. (2, 1) C. (1, -2) D. (-2, 1) Solve the system of equations. 3x - 4y = -10 and 5x + 8y = -2 A. (1, 2) B. (2, 1) C. (1, -2) D. (-2, 1) @Mathematics

OpenStudy (hoblos):

D

OpenStudy (anonymous):

3x-4y=-10 3x=-10+4y x=(-10/3)+(4/3)y Plug into second equation and solve for y: 5x+8y=-2 5[(-10/3)+(4/3)y]+8y=-2 (-50/3)+(20/3)y+8y=-2 (44/3)y=(44/3) y=1 Plug y back into first equation and solve for x: 3x-4y=-10 x=(-10/3)+(4/3)y x=(-10/3)+(4/3)(1) x=-2 Final answer: x=-2 y=1